OpenStudy (anonymous):

49. The depth of water behind the Hoover Dam is 220 m. Show that the water pressure at the base of this dam is 2160 kPa.

OpenStudy (anonymous):

P=p*g*h P=1000*9.8*220 =2156000 2156000/1000= 2160kpa

OpenStudy (anonymous):

Medal please